### v1

Abstract
chiappori2025changes study several indices that measure the assortativeness of a matching, including the aggregate likelihood ratio. We provide a counterexample showing that their axiomatization of the aggregate likelihood ratio is not valid as stated. We identify the exact class of indices characterized by the axioms in chiappori2025changes. We then show that the axiomatization of the aggregate likelihood ratio can be recovered by adding new axioms.
1 Introduction
The degree of assortative matching between men and women—how similar partners are in terms of education, age, income, and other socioeconomic characteristics—has long been viewed as a central determinant of inequality and welfare in modern economies (chiappori2023mating). If high-income or highly educated individuals increasingly marry one another, household income inequality may rise even when individual wage distributions remain stable. Understanding the extent and evolution of assortative matching is therefore crucial for interpreting long-run trends in inequality and social mobility.
In a pioneering study, chiappori2025changes propose an axiomatic foundation for indices of assortative matching and provide empirical evidence on changes in marital sorting in the United States. Their framework offers a valuable basis for interpreting trends in marital sorting in subsequent work. In this note, we show that the axiomatization of one measure, the aggregate likelihood ratio, is not correct as currently stated.111eika2019educational define a type-specific likelihood ratio as the ratio of the observed probability that a man and a woman of the same type (e.g., age) marry to the corresponding probability under random matching. The aggregate likelihood ratio is the weighted average of these type-specific likelihood ratios, with weights given by the expected mass of like-type couples under random matching. The measure has been applied to sociology and demographic studies such as pesando2021educational. We identify the exact class of indices characterized by the axioms in chiappori2025changes (Theorem 1). We then propose a refinement that corrects the statement and restores the intended result, while preserving the spirit and practical usefulness of their original approach (Theorem 2).
2 Model
Our notation follows that of chiappori2025changes. We consider two-type markets, i.e., there are two types of men and two types of women. Let $\mathcal{M}=\mathbb{R}_{\geq 0}^{4}\backslash\{\mathbf{0}\}$ denote the entire collection of possible matching patterns. We represent a matching pattern $M\in\mathcal{M}$ by a $2\times 2$ matrix given by
[…]
If the two types are “high income” and “low income”, then $a$ represents the number of matches between high-income men and high-income women, $b$ represents the number of matches between high-income men and low-income women, $c$ represents the number of matches between low-income men and high-income women, and $d$ represents the number of matches between low-income men and low-income women.
We call $M\in\mathcal{M}$ a matching matrix, or simply a matching. We often represent a matching in the vector form as $M=(a,b,c,d)$. For each $M\in\mathcal{M}$, let $|M|:=a+b+c+d$ denote the population size. We write $M>\mathbf{0}$ to mean that every entry of $M$ is positive. Such $M$ is called a positive matching matrix. Let $\mathcal{M}_{>0}$ denote the set of all positive matching matrices.
An index is a function $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. For each $M\in\mathcal{M}$, $I(M)$ represents how assortative the matching matrix is. Among several indices considered in chiappori2025changes, we focus on the following one.
  • The aggregate likelihood ratio (abbreviated ALR) is defined by
    $I_{L}(M)=\frac{a+d}{|M|}/\left(\frac{a+b}{|M|}\frac{a+c}{|M|}+\frac{d+b}{|M|}\frac{d+c}{|M|}\right)\text{ for all }M\in\mathcal{M}.$
    Equivalently,
    $I_{L}(M)=(a+d)/\left(\frac{(a+b)(a+c)}{|M|}+\frac{(d+b)(d+c)}{|M|}\right)\text{ for all }M\in\mathcal{M}.$
Notice that ALR is not well-defined for matching matrices $M=(a,b,c,d)$ such that $(a+b)(a+c)+(d+b)(d+c)=0$ (this equation holds if $a=b=d=0$ or $a=c=d=0$). Therefore, in the following analysis, we confine our attention to the class of matching matrices $\tilde{\mathcal{M}}\subseteq\mathcal{M}$ for which ALR is well-defined.
$\tilde{\mathcal{M}}=\{(a,b,c,d)\in\mathcal{M}\mid(a+b)(a+c)+(d+b)(d+c)\neq 0\}.$
We revisit the axioms in chiappori2025changes. Consider a domain $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ with $\mathcal{M}^{\prime}\neq\emptyset$. We define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$. [^2] We note that chiappori2025changes’s (chiappori2025changes) define axioms for an index $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. The following three axioms are called “invariance axioms”.
  • Scale Invariance. For every $M\in\mathcal{M}^{\prime}$ and $\lambda>0$ with $\lambda M\in\mathcal{M}^{\prime}$, we have $I(M)=I(\lambda\cdot M)$.
  • Side Invariance.
[…]
  • Type Invariance.
[…]
Monotonicity axioms state that if the underlying matching matrix changes in a certain way, then the assortativeness increases.
  • Marginal Monotonicity. For every pair of positive matching matrices $M=(a,b,c,d)\in\mathcal{M}^{\prime}$ and $M^{\prime}=\left(a^{\prime},b^{\prime},c^{\prime},d^{\prime}\right)\in\mathcal{M}^{\prime}$ with the same marginal distributions (i.e., $a+c=a^{\prime}+c^{\prime},a+b=a^{\prime}+b^{\prime},d+b=d^{\prime}+b^{\prime}$, and $\left.d+c=d^{\prime}+c^{\prime}\right)$, $I(M)>I(M^{\prime})$ if and only if $a>a^{\prime}$ (equivalently, $b<b^{\prime},c<c^{\prime}$, or $\left.d>d^{\prime}\right)$.
To axiomatize ALR, chiappori2025changes introduced the following axiom.
  • Random Decomposability. For every pair of positive matching matrices $M,M^{\prime}\in\mathcal{M}^{\prime}$,
    $I\left(M+M^{\prime}\right)=\frac{r(M)}{r\left(M+M^{\prime}\right)}I(M)+\frac{r\left(M^{\prime}\right)}{r\left(M+M^{\prime}\right)}I\left(M^{\prime}\right),$
    where, for each $M\in\mathcal{M}^{\prime}$,
    $r(M)$ $:=\frac{a+b}{|M|}\frac{a+c}{|M|}|M|+\frac{d+b}{|M|}\frac{d+c}{|M|}|M|$
    $=\frac{(a+b)(a+c)+(d+b)(d+c)}{|M|}.$ (1)
We reproduce Theorem 3 of chiappori2025changes.
Theorem CDMZ3.
An index satisfies the three invariance axioms, marginal monotonicity, and random decomposability if and only if it is a positive multiple of aggregate likelihood ratio.
3 Counterexample to Theorem CDMZ3
To see that this theorem does not hold for an index $I:\hat{\mathcal{M}}\rightarrow\mathbb{R}_{\geq 0}$, consider the following variant of $I_{L}$.
$I^{\prime}_{L}(M)=\Bigl(a+d+\frac{b}{2}+\frac{c}{2}\Bigr)/\left(\frac{(a+b)(a+c)}{|M|}+\frac{(d+b)(d+c)}{|M|}\right)\text{ for all }M\in\tilde{\mathcal{M}}$.
In the next section we prove that $I^{\prime}_{L}$ satisfies all the axioms in Theorem CDMZ3; see the proof of the “if” direction of Theorem 1. We show that $I^{\prime}_{L}$ is not a positive multiple of $I_{L}$. Suppose, for contradiction, that $I^{\prime}_{L}$ is a positive multiple of $I_{L}$. Then, for every $M,M^{\prime}\in\tilde{\mathcal{M}}$,
$I_{L}(M)>I_{L}(M^{\prime})\Longleftrightarrow I^{\prime}_{L}(M)>I^{\prime}_{L}(M^{\prime}).$ (2)
Consider the following two matrices.
[…]
Then,
$I_{L}(M)=1>\frac{21}{23}=I_{L}\left(M^{\prime}\right),$
$I_{L}^{\prime}(M)=\frac{3}{2}<\frac{35}{23}=I_{L}^{\prime}\left(M^{\prime}\right).$
We obtain a contradiction to (2).
4 Recovering axiomatization of the aggregate likelihood ratio
In this section, we recover the axiomatization of ALR in chiappori2025changes’s (chiappori2025changes). To this end, we first identify the exact class of indices that satisfy the axioms in Theorem CDMZ3.
Theorem 1.
An index $I:\tilde{\mathcal{M}}\rightarrow\mathbb{R}_{\geq 0}$ satisfies the three invariance axioms, marginal monotonicity, and random decomposability if and only if the following two conditions hold:
  • (i)
    On the class of positive matching matrices $\mathcal{M}_{>0}$, there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that,
    $I\Bigl(\begin{tabular}[]{c|c}$a$&$b$\\ \hline\cr$c$&$d$\end{tabular}\Bigr)=\cfrac{\alpha a+\beta b+\beta c+\alpha d}{r(M)}\text{ for all }M=\Bigl(\begin{tabular}[]{c|c}$a$&$b$\\ \hline\cr$c$&$d$\end{tabular}\Bigr)\in\mathcal{M}_{>0}.$ (7)
  • (ii)
    On the class of matching matrices $\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, $I$ is an index that satisfies the three invariance axioms.
Proof.
Proof of the “if” direction: It is clear that $I$ satisfies all the axioms on $\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, because on this domain, marginal monotonicity and random decomposability have no bite (recall that these axioms are defined only for two positive matching matrices in $\mathcal{M}_{>0}$).
We show that $I$ satisfies all the axioms on the domain $\mathcal{M}_{>0}$. Suppose that there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that $I$ is given as in (7). This index satisfies side invariance because it is symmetric with respect to $b$ and $c$. Similarly, it satisfies type invariance because it is symmetric with respect to $a$ and $d$. Furthermore, it satisfies scale invariance because, after multiplying a matching matrix $M\in\tilde{\mathcal{M}}$ by $\lambda>0$, both the numerator and the denominator of $I(M)$ are multiplied by $\lambda$. In the following we show that $I$ satisfies marginal monotonicity and random decomposability.
Proof that $I$ satisfies marginal monotonicity: Take $\epsilon\in\mathbb{R}$ and consider two positive matching matrices $M,M_{\epsilon}\in\mathcal{M}_{>0}$ with the same marginal distributions, i.e.,
[…]
Since
$\left|M_{\varepsilon}\right|=(a+\varepsilon)+(b-\varepsilon)+(c-\varepsilon)+(d+\varepsilon)=|M|,$
the population size is the same between the two matching matrices. The following equations also hold:
$\begin{array}[]{ll}(a+\varepsilon)+(b-\varepsilon)=a+b,&(a+\varepsilon)+(c-\varepsilon)=a+c,\\ (d+\varepsilon)+(b-\varepsilon)=d+b,&(d+\varepsilon)+(c-\varepsilon)=d+c.\end{array}$
Therefore, $r(M)=r(M_{\epsilon})$. Moreover,
Numerator of $I(M_{\epsilon})$ $=\alpha\cdot(a+\epsilon)+\alpha\cdot(d+\epsilon)+\beta\cdot(b-\epsilon)+\beta\cdot(c-\epsilon)$
$=\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c+2\epsilon\cdot(\alpha-\beta)$
Numerator of $I(M)$ $=\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c.$
Under $\alpha>\beta$, we have
$I(M_{\epsilon})>I(M)\>\Longleftrightarrow\>\epsilon>0,$
as desired.
Proof that $I$ satisfies random decomposability: For each $M=(a,b,c,d)\in\mathcal{M}_{>0}$, we define
$A(M):=\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c.$
Then
$I(M)=\frac{A(M)}{r(M)}.$
Consider $M,M^{\prime}\in\mathcal{M}_{>0}$ and let $M^{\prime\prime}:=M+M^{\prime}$. Since $A(\cdot)$ is linear in $(a,b,c,d)$,
$A(M^{\prime\prime})=A(M)+A\left(M^{\prime}\right).$
Then,
$\frac{r(M)}{r(M^{\prime\prime})}I(M)+\frac{r\left(M^{\prime}\right)}{r(M^{\prime\prime})}I\left(M^{\prime}\right)=\frac{r(M)}{r(M^{\prime\prime})}\frac{A(M)}{r(M)}+\frac{r\left(M^{\prime}\right)}{r(M^{\prime\prime})}\frac{A\left(M^{\prime}\right)}{r\left(M^{\prime}\right)}=\frac{A(M)+A\left(M^{\prime}\right)}{r(M^{\prime\prime})}=\frac{A(M^{\prime\prime})}{r(M^{\prime\prime})}=I(M^{\prime\prime}),$
which establishes random decomposability.
Proof of the “only if” direction: Let $I$ be an index that satisfies the three invariance axioms, marginal monotonicity, and random decomposability. It is clear that, on the domain of matching matrices $\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, $I$ is an index that satisfies the three invariance axioms.
It suffices to prove that, on the domain $\mathcal{M}_{>0}$, $I$ is given as stated. We define $\tilde{I}:\mathcal{M}_{>0}\rightarrow\mathbb{R}_{\geq 0}$ by
$\tilde{I}(M)=r(M)\cdot I(M)\text{ for all }M\in\mathcal{M}_{>0}.$ (8)
For each $M\in\mathcal{M}_{>0}$, since $r(M)\geq 0$ and $I(M)\geq 0$, we have $\tilde{I}(M)\geq 0$.
For every pair of positive matching matrices $M,M^{\prime}\in\mathcal{M}_{>0}$, it holds that
$\tilde{I}(M+M^{\prime})$ $=r(M+M^{\prime})\cdot I(M+M^{\prime})$
$=r(M)\cdot I(M)+r(M^{\prime})\cdot I(M^{\prime})$
$=\tilde{I}(M)+\tilde{I}(M^{\prime}),$
where the second equality follows from random decomposability. Therefore, $\tilde{I}$ is additive.
For $M\in\mathcal{M}_{>0}$ and $\lambda>0$,
$\tilde{I}(\lambda M)$ $=r(\lambda M)\cdot I(\lambda M)$
$=\lambda\cdot r(M)\cdot I(\lambda M)$
$=\lambda\cdot r(M)\cdot I(M)$
$=\lambda\cdot\tilde{I}(M),$
where the second equality follows from the definition of $r(M)$ (recall (1)) and the third equality follows from scale invariance. Therefore, $\tilde{I}$ satisfies positive homogeneity of degree 1.
Let
[…]
Also let
$\alpha^{\prime}:=\tilde{I}(M_{1}),\>\>\beta^{\prime}:=\tilde{I}(M_{2}),\>\>\gamma^{\prime}:=\tilde{I}(M_{3}),\>\>\delta^{\prime}:=\tilde{I}(M_{4}).$ (9)
By side invariance and type invariance,
$\beta^{\prime}=\gamma^{\prime},\>\>\alpha^{\prime}=\delta^{\prime}.$ (10)
Consider $M\in\mathcal{M}_{>0}$. Since $M_{1},M_{2},M_{3},M_{4}$ form a basis of $\mathbb{R}^{4}$, there exist coefficients $s_{1},s_{2},s_{3},s_{4}$ such that
$M=s_{1}M_{1}+s_{2}M_{2}+s_{3}M_{3}+s_{4}M_{4}.$
The coefficients are explicitly given by
$s_{1}=a-\frac{a+b+c+d}{5},$
$s_{2}=b-\frac{a+b+c+d}{5},$
$s_{3}=c-\frac{a+b+c+d}{5},$
$s_{4}=d-\frac{a+b+c+d}{5}.$ (11)
Define
$t_{i}^{\prime}:=\max\left\{s_{i},0\right\},\quad t_{i}:=\max\left\{-s_{i},0\right\}\quad(i=1,2,3,4)$
For each $i=1,\dots,4$, we have $t_{i},t_{i}^{\prime}\geq 0$ and $t_{i}^{\prime}-t_{i}=s_{i}$. By the latter equation,
$M=\sum_{i=1}^{4}\left(t_{i}^{\prime}-t_{i}\right)M_{i}\quad\Longleftrightarrow\quad M+\sum_{i=1}^{4}t_{i}M_{i}=\sum_{i=1}^{4}t_{i}^{\prime}M_{i}.$ (12)
Since $\tilde{I}$ satisfies additivity and positive homogeneity of degree 1, [^3] We note that the equation below holds including the case that $t_{i}=0$ for some $i=1,\dots,4$.
$\tilde{I}\Bigl(M+\sum_{i=1}^{4}t_{i}M_{i}\Bigr)=\tilde{I}(M)+\sum_{i=1}^{4}t_{i}\tilde{I}(M_{i}).$
Similarly,
$\tilde{I}\Bigl(\sum_{i=1}^{4}t_{i}^{\prime}M_{i}\Bigr)=\sum_{i=1}^{4}t^{\prime}_{i}\tilde{I}(M_{i}).$
The above two displayed equations together with (12) imply
$\tilde{I}(M)=\sum_{i=1}^{4}t^{\prime}_{i}\tilde{I}(M_{i})-\sum_{i=1}^{4}t_{i}\tilde{I}(M_{i})=\sum_{i=1}^{4}s_{i}\tilde{I}(M_{i}).$
By substituting (9), (10), and (11) into the above equation,
$\tilde{I}(M)=\frac{3\alpha^{\prime}-2\beta^{\prime}}{5}a+\frac{-2\alpha^{\prime}+3\beta^{\prime}}{5}b+\frac{-2\alpha^{\prime}+3\beta^{\prime}}{5}c+\frac{3\alpha^{\prime}-2\beta^{\prime}}{5}d.$
By setting
$\alpha:=\frac{3\alpha^{\prime}-2\beta^{\prime}}{5},\;\>\beta:=\frac{-2\alpha^{\prime}+3\beta^{\prime}}{5},$
we obtain
$\tilde{I}(M)=\alpha a+\beta b+\beta c+\alpha d.$
Since $\tilde{I}(M)=r(M)\cdot I(M)$, $I$ is given in the form of (7).
As noted after (8), $\tilde{I}(M)\geq 0$ for all $M\in\tilde{\mathcal{M}}$, which implies $\alpha\geq 0$ and $\beta\geq 0$ hold. Indeed, if $\alpha<0$, by letting $a,d\rightarrow+\infty$ and $b,c\rightarrow 0$, we obtain a matching matrix $M=(a,b,c,d)\in\mathcal{M}_{>0}$ for which $\tilde{I}(M)<0$, a contradiction to $\tilde{I}(M)\geq 0$. Similarly, $\beta<0$ yields a contradiction. It remains to prove that $\alpha>\beta$. Take $\epsilon\in\mathbb{R}$ with $\epsilon>0$. Consider two positive matching matrices $M,M_{\epsilon}\in\mathcal{M}_{>0}$ such that
[…]
As shown in the proof of the “if” part, we have $r(M)=r(M_{\epsilon})$. By marginal monotonicity of $I$,
$\alpha\cdot(a+\epsilon)+\alpha\cdot(d+\epsilon)+\beta\cdot(b-\epsilon)+\beta\cdot(c-\epsilon)>\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c.$
This inequality implies $2\epsilon\cdot(\alpha-\beta)>0$. Since $\epsilon>0$, we obtain $\alpha>\beta$, as desired. ∎
Theorem 1 suggests that, to pin down the class of positive multiples of ALR, we need additional axioms that force $\beta=0$ in (7). We consider an axiom that is interpreted as the opposite of the maximum homogamy axiom in chiappori2025changes (defined in Section 5). As in Section 2, we define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}^{\prime}\subseteq\mathcal{M}$, $\mathcal{M}^{\prime}\neq\emptyset$.
  • Maximum Heterogamy. For each $M\in\mathcal{M}^{\prime}$ and $(a,b,c,d)\in\mathcal{M}^{\prime}$ such that $b>0,c>0,a=d=0$, we have $I(M)\geq I(a,b,c,d)$.
We also introduce the continuity axiom, which states that if the underlying matching matrix changes slightly, then the index changes slightly as well.
  • Continuity: $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$ is a continuous function.
Theorem 2.
An index $I:\tilde{\mathcal{M}}\rightarrow\mathbb{R}_{\geq 0}$ satisfies the three invariance axioms, marginal monotonicity, random decomposability, maximum heterogamy, and continuity if and only if there exists $\alpha\in\mathbb{R}$ with $\alpha>0$ such that
[…]
Proof.
Proof of the “if” direction: Suppose that there exists $\alpha>0$ such that $I$ is given as in (17). By following the same argument as in the “if” direction of Theorem 1, we can show that $I$ satisfies the three invariance axioms, marginal monotonicity, and random decomposability. Since $I$ is a linear function, it satisfies continuity. For each $(a,b,c,d)\in\tilde{\mathcal{M}}$ with $b>0,c>0,a=d=0$, we have $I(a,b,c,d)=0$. Therefore, maximum heterogamy holds.
Proof of the “only if” direction: By Theorem 1, there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that $I$ is given as in (7) on the domain $\mathcal{M}_{>0}$. If $I$ satisfies maximum heterogamy, then $\beta=0$. Indeed, if $\beta>0$, by letting $b,c\rightarrow+\infty$ while keeping $a=d=0$, the value of $I(a,b,c,d)$ grows arbitrarily large, contradicting maximum heterogamy.
Note that for any $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, there exists a sequence $\{M^{k}\}_{k=1}^{\infty}\subseteq\mathcal{M}_{>0}$ such that $M^{k}\rightarrow M$. Therefore, by continuity, the value of $I(M)$ for $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$ is also given by (17). ∎
5 Concluding remarks
We have shown that Theorem 3 of chiappori2025changes does not hold as currently stated and we proposed a modification. We note that Theorem 2 of chiappori2025changes does not hold as currently stated either. Specifically, chiappori2025changes consider the following index called the normalized trace:
$I_{tr}(M)=\left\{\begin{array}[]{cl}1&\text{ if }bc=0,\\ \frac{a+d}{a+b+c+d}\in(0,1)&\text{ if }abcd\neq 0,\\ 0&\text{ if }ad=0.\end{array}\right$.
We remark that this index is not well-defined on the entire domain of matchings $\mathcal{M}$; the conditions “bc=0bc=0” (first line) and “ad=0ad=0” (third line) overlap, so a matching matrix such as $(1,1,0,0)$ satisfies both clauses and is assigned the values 1 and 0 simultaneously. Therefore, we restrict the domain of an index to the set of matchings $\hat{\mathcal{M}}$ defined by
$\hat{\mathcal{M}}=\{M\in\mathcal{M}\mid bc\neq 0\text{ or }ad\neq 0\}.$
The following is Theorem 2 of chiappori2025changes.
Theorem CDMZ2.
The normalized trace is the unique index, up to positive affine transformation, that satisfies the three invariance axioms, marginal monotonicity, maximum homogamy, and population decomposability.
The last two axioms in the statement are defined as follows:
  • Maximum Homogamy. For each $M\in\hat{\mathcal{M}}$ and $(a,b,c,d)\in\hat{\mathcal{M}}$ such that $ad>0$ and $bc=0$, we have $I(a,b,c,d)\geq I(M)$.
  • Population Decomposability. For each pair of positive matchings $M,M^{\prime}\in\hat{\mathcal{M}}$,
    $I\left(M+M^{\prime}\right)=\frac{|M|}{|M|+\left|M^{\prime}\right|}I(M)+\frac{\left|M^{\prime}\right|}{|M|+\left|M^{\prime}\right|}I\left(M^{\prime}\right).$
To see that Theorem CDMZ2 does not hold, consider the following index defined on $\hat{\mathcal{M}}$.
$I^{\prime}_{tr}(M)=\left\{\begin{array}[]{cl}1&\text{ if }bc=0,\\ \frac{a+d+\frac{b}{2}+\frac{c}{2}}{a+b+c+d}\in(0,1)&\text{ if }abcd\neq 0,\\ 0&\text{ if }ad=0.\end{array}\right$.
One can verify that this modified index satisfies the three invariance axioms, marginal monotonicity, maximum homogamy, and population decomposability. [^4] This modified index also satisfies other two monotonicity axioms in chiappori2025changes, namely diagonal monotonicity and off-diagonal monotonicity. However, this modified index is not a positive affine transformation of the normalized trace. To see this, let $M=(a,b,c,d)\in\hat{\mathcal{M}}$. We consider two cases.
Case 1: Suppose $bc=0,ad>0$. Then,
$I_{tr}(M)=1,\quad I_{tr}^{\prime}(M)=1$
Hence
$1=\alpha\cdot 1+\beta\quad\Rightarrow\quad\alpha+\beta=1.$
Case 2: Suppose $ad=0,bc>0$. Then
$I_{tr}(M)=0,\quad I_{tr}^{\prime}(M)=0$
Hence
$0=\alpha\cdot 0+\beta\quad\Rightarrow\quad\beta=0.$
Together with $\alpha+\beta=1$, we have $\alpha=1$. This means that
$I_{tr}(M)=I^{\prime}_{tr}(M)\text{ for all }M\in\mathcal{M}.$ (18)
Let $M=(1,1,1,1)$. Then,
$I_{tr}(M)$ $=\frac{1}{2},$
$I_{tr}^{\prime}(M)$ $=\frac{3}{4}$
We obtain a contradiction to (18).
By adopting a technique similar to that developed in Section 4, we can recover the axiomatization of the normalized trace. We will provide details in an updated version of this note.

### v2

@@
-
+This version: March 12, 2026 )
 Abstract
-chiappori2025changes study several indices that measure the assortativeness of a matching, including the aggregate likelihood ratio.
+Chiappori et al.
+(2025) study several indices of assortativeness in matching, including the aggregate likelihood ratio and the odds ratio.
 We provide a counterexample showing that their axiomatization of the aggregate likelihood ratio is not valid as stated.
-We identify the exact class of indices characterized by the axioms in chiappori2025changes.
+We identify the exact class of indices characterized by the axioms in Chiappori et al.
+(2025).
 We then show that the axiomatization of the aggregate likelihood ratio can be recovered by adding new axioms.
+In addition, we point out errors in the axiomatizations of other measures in Chiappori et al.
+(2025).
+Finally, we offer a generalization of the odds ratio from two-type markets to multi-type markets.
 1 Introduction
-The degree of assortative matching between men and women—how similar partners are in terms of education, age, income, and other socioeconomic characteristics—has long been viewed as a central determinant of inequality and welfare in modern economies (chiappori2023mating).
+The degree of assortative matching between men and women—how similar partners are in terms of education, age, income, and other socioeconomic characteristics—has long been viewed as a central determinant of inequality and welfare in modern economies (Chiappori and Salanié 2023).
 If high-income or highly educated individuals increasingly marry one another, household income inequality may rise even when individual wage distributions remain stable.
 Understanding the extent and evolution of assortative matching is therefore crucial for interpreting long-run trends in inequality and social mobility.
-In a pioneering study, chiappori2025changes propose an axiomatic foundation for indices of assortative matching and provide empirical evidence on changes in marital sorting in the United States.
+In a pioneering study, Chiappori et al.
+(2025) propose an axiomatic foundation for indices of assortative matching and provide empirical evidence on changes in marital sorting in the United States.
 Their framework offers a valuable basis for interpreting trends in marital sorting in subsequent work.
-In this note, we show that the axiomatization of one measure, the aggregate likelihood ratio, is not correct as currently stated.111eika2019educational define a type-specific likelihood ratio as the ratio of the observed probability that a man and a woman of the same type (e.g., age) marry to the corresponding probability under random matching.
+In this note, we show that the axiomatization of one measure, the aggregate likelihood ratio, is not correct as currently stated.
+[^1] Eika et al.
+(2019) define a type-specific likelihood ratio as the ratio of the observed probability that a man and a woman of the same type (e.g., age) marry to the corresponding probability under random matching.
 The aggregate likelihood ratio is the weighted average of these type-specific likelihood ratios, with weights given by the expected mass of like-type couples under random matching.
-The measure has been applied to sociology and demographic studies such as pesando2021educational.
-We identify the exact class of indices characterized by the axioms in chiappori2025changes (Theorem 1).
+The measure has been applied to sociology and demographic studies such as Pesando (2021).
+We identify the exact class of indices characterized by the axioms in Chiappori et al.
+(2025) (Theorem 1).
 We then propose a refinement that corrects the statement and restores the intended result, while preserving the spirit and practical usefulness of their original approach (Theorem 2).
+Chiappori et al.
+(2025) axiomatize two measures other than the aggregate likelihood ratio: the odds ratio and the normalized trace.
+We show that the axiomatizations of these masures are not correct either.
+In the final section, we discuss a generalization of the odds ratio from two-type markets to multi-type markets.
 2 Model
-Our notation follows that of chiappori2025changes.
+Our notation follows that of Chiappori et al.
+(2025).
 We consider two-type markets, i.e., there are two types of men and two types of women.
@@
 For each $M\in\mathcal{M}$, $I(M)$ represents how assortative the matching matrix is.
-Among several indices considered in chiappori2025changes, we focus on the following one.
+Among several indices considered in Chiappori et al.
+(2025), we focus on the following one.
   • The aggregate likelihood ratio (abbreviated ALR) is defined by
@@
 $\tilde{\mathcal{M}}=\{(a,b,c,d)\in\mathcal{M}\mid(a+b)(a+c)+(d+b)(d+c)\neq 0\}.$
-We revisit the axioms in chiappori2025changes.
+We revisit the axioms in Chiappori et al.
+(2025).
 Consider a domain $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ with $\mathcal{M}^{\prime}\neq\emptyset$.
 We define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$.
-[^2] We note that chiappori2025changes’s (chiappori2025changes) define axioms for an index $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$.
+[^2] We note that Chiappori et al.’s (2025) define axioms for an index $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$.
 The following three axioms are called “invariance axioms”.
@@
 For every pair of positive matching matrices $M=(a,b,c,d)\in\mathcal{M}^{\prime}$ and $M^{\prime}=\left(a^{\prime},b^{\prime},c^{\prime},d^{\prime}\right)\in\mathcal{M}^{\prime}$ with the same marginal distributions (i.e., $a+c=a^{\prime}+c^{\prime},a+b=a^{\prime}+b^{\prime},d+b=d^{\prime}+b^{\prime}$, and $\left.d+c=d^{\prime}+c^{\prime}\right)$, $I(M)>I(M^{\prime})$ if and only if $a>a^{\prime}$ (equivalently, $b<b^{\prime},c<c^{\prime}$, or $\left.d>d^{\prime}\right)$.
-To axiomatize ALR, chiappori2025changes introduced the following axiom.
+To axiomatize ALR, Chiappori et al. (2025) introduced the following axiom.
   • Random Decomposability. For every pair of positive matching matrices $M,M^{\prime}\in\mathcal{M}^{\prime}$,
@@
     $=\frac{(a+b)(a+c)+(d+b)(d+c)}{|M|}.$ (1)
-We reproduce Theorem 3 of chiappori2025changes.
+We reproduce Theorem 3 of Chiappori et al. (2025).
 Theorem CDMZ3.
@@
 4 Recovering axiomatization of the aggregate likelihood ratio
-In this section, we recover the axiomatization of ALR in chiappori2025changes’s (chiappori2025changes).
-To this end, we first identify the exact class of indices that satisfy the axioms in Theorem CDMZ3.
+In this section, we recover the axiomatization of ALR in Chiappori et al. (2025). To this end, we first identify the exact class of indices that satisfy the axioms in Theorem CDMZ3.
 Theorem 1.
@@
 Theorem 1 suggests that, to pin down the class of positive multiples of ALR, we need additional axioms that force $\beta=0$ in (7).
-We consider an axiom that is interpreted as the opposite of the maximum homogamy axiom in chiappori2025changes (defined in Section 5).
+We consider an axiom that is interpreted as the opposite of the maximum homogamy axiom in Chiappori et al.
+(2025) (defined in Section 5).
 As in Section 2, we define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}^{\prime}\subseteq\mathcal{M}$, $\mathcal{M}^{\prime}\neq\emptyset$.
-  • Maximum Heterogamy. For each $M\in\mathcal{M}^{\prime}$ and $(a,b,c,d)\in\mathcal{M}^{\prime}$ such that $b>0,c>0,a=d=0$, we have $I(M)\geq I(a,b,c,d)$.
+  • Maximum Heterogamy. For each $M\in\mathcal{M}_{>0}$ and $(a,b,c,d)\in\mathcal{M}^{\prime}$ such that $b>0,c>0,a=d=0$, we have $I(M)\geq I(a,b,c,d)$.
 We also introduce the continuity axiom, which states that if the underlying matching matrix changes slightly, then the index changes slightly as well.
@@
 Proof of the “only if” direction: By Theorem 1, there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that $I$ is given as in (7) on the domain $\mathcal{M}_{>0}$.
+Note that for any $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, there exists a sequence $\{M^{k}\}_{k=1}^{\infty}\subseteq\mathcal{M}_{>0}$ such that $M^{k}\rightarrow M$.
+Therefore, by continuity, the value of $I(M)$ for $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$ is also given by (17).
 If $I$ satisfies maximum heterogamy, then $\beta=0$.
 Indeed, if $\beta>0$, by letting $b,c\rightarrow+\infty$ while keeping $a=d=0$, the value of $I(a,b,c,d)$ grows arbitrarily large, contradicting maximum heterogamy.
-Note that for any $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, there exists a sequence $\{M^{k}\}_{k=1}^{\infty}\subseteq\mathcal{M}_{>0}$ such that $M^{k}\rightarrow M$.
-Therefore, by continuity, the value of $I(M)$ for $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$ is also given by (17).
 ∎
-5 Concluding remarks
-We have shown that Theorem 3 of chiappori2025changes does not hold as currently stated and we proposed a modification.
-We note that Theorem 2 of chiappori2025changes does not hold as currently stated either.
-Specifically, chiappori2025changes consider the following index called the normalized trace:
+5 Remarks on axiomatizations of other measures
+We have shown that Theorem 3 of Chiappori et al.
+(2025) does not hold as currently stated and we proposed a modification.
+We note that Theorems 1 and 2 of Chiappori et al.
+(2025) do not hold as currently stated either.
+5.1 Odds ratio
+For $(a,b,c,d)\in\mathcal{M}$, we define the odds ratio by
+$I_{O}(M)=\left\{\begin{array}[]{l}\frac{ad}{bc}\text{ if }b\neq 0\text{ and }c\neq 0,\\ +\infty\text{ if }bc=0.\end{array}\right$.
+A preorder $\succeq$ is a reflexive and transitive binary relation on $\mathcal{M}$. We revisit Chiappori et al.’s (2025) axioms on a preorder $\succeq$.
+  • Marginal Independence. For any $M\in\mathcal{M}$ and $\lambda>0$,
+    $a$ $b$ $c$ $d$ $\sim$ $\lambda a$ $\lambda b$ $c$ $d$ $\sim$ $a$ $b$ $\lambda c$ $\lambda d$ $\sim$ $\lambda a$ $b$ $\lambda c$ $d$ $\sim$ $a$ $\lambda b$ $c$ $\lambda d$ .
+  • Maximum Homogamy. For any $M\in\mathcal{M},ad>0$, and $bc=0,(a,b,c,d)\succeq M$.
+  • Marginal Monotonicity (for preorders).
+For every pair of positive matching matrices $M=(a,b,c,d)\in\mathcal{M}^{\prime}$ and $M^{\prime}=\left(a^{\prime},b^{\prime},c^{\prime},d^{\prime}\right)$ with the same marginal distributions (i.e., $a+c=a^{\prime}+c^{\prime},a+b=a^{\prime}+b^{\prime},d+b=d^{\prime}+b^{\prime}$, and $\left.d+c=d^{\prime}+c^{\prime}\right)$, $M\succ M^{\prime}$ if and only if $a>a^{\prime}$ (equivalently, $b<b^{\prime},c<c^{\prime}$, or $\left.d>d^{\prime}\right)$.
+The following is Theorem 1 of Chiappori et al. (2025).
+Theorem CDMZ1.
+The odds ratio induces the unique preorder that satisfies marginal monotonicity, marginal independence, and maximum homogamy.
+To introduce a counterexample to this theorem, we define three subsets of $\mathcal{M}$.
+  • $\mathcal{M}_{\infty}=\{(a,b,c,d)\mid bc=0,ad>0\}$.
+  • $\widetilde{\mathcal{M}}_{\infty}=\{(a,b,c,d)\mid bc=0,ad=0\}$
+  • $\mathcal{M}_{0}=\{(a,b,c,d)\mid bc>0\text{ and }ad=0\}$.
+Then, it holds that $\mathcal{M}=\mathcal{M}_{\infty}\cup\widetilde{\mathcal{M}}_{\infty}\cup\mathcal{M}_{0}\cup\mathcal{M}_{>0}$.
+Note that $\mathcal{M}_{\infty}$ and $\widetilde{\mathcal{M}}_{\infty}$ denote the set of matrices to which the odds ratio assigns $\infty$, $\mathcal{M}_{0}$ denotes the set of matrices to which the odds ratio assigns 0, and $\mathcal{M}_{>0}$ denotes the set of matrices to which the odds ratio assigns positive real number.
+We define a preorder $\succeq^{*}$ as follows:
+  • (1)
+    for each $M\in\mathcal{M}_{\infty}$ and $M^{\prime}\in\widetilde{\mathcal{M}}_{\infty}\cup\mathcal{M}_{0}\cup\mathcal{M}_{>0}$, we have $M\succ^{*}M^{\prime}$.
+    for each $M,M^{\prime}\in\mathcal{M}_{\infty}$, we have $M\sim^{*}M^{\prime}$.
+  • (2)
+    for each $M\in\widetilde{\mathcal{M}}_{\infty}$ and $M^{\prime}\in\mathcal{M}_{0}\cup\mathcal{M}_{>0}$, we have $M\succ^{*}M^{\prime}$.
+    for each $M,M^{\prime}\in\widetilde{\mathcal{M}}_{\infty}$, we have $M\sim^{*}M^{\prime}$.
+  • (3)
+    for each $M\in\mathcal{M}_{0}$ and $M^{\prime}\in\mathcal{M}_{>0}$, we have $M\succ^{*}M^{\prime}$.
+    for each $M,M^{\prime}\in\mathcal{M}_{0}$, we have $M\sim^{*}M^{\prime}$.
+  • (4)
+    for each $M,M^{\prime}\in\mathcal{M}_{>0}$, we have $M\succeq^{*}M^{\prime}$ if and only if $I_{O}(M)\geq I_{O}(M^{\prime})$.
+Then, $\succeq^{*}$ satisfies all the three axioms of Theorem CDMZ1.
+First, $\succeq^{*}$ satisfies marginal monotonicity because this axiom bears only on matrices in $\mathcal{M}_{>0}$, and in this space, $\succeq^{*}$ is induced from the odds ratio.
+Moreover, $\succeq^{*}$ satisfies maximum homogamy by Condition (1).
+We show that $\succeq^{*}$ satisfies marginal independence.
+  • For $M\in\mathcal{M}_{\infty}$ and $\lambda>0$, we have $\lambda M\in\mathcal{M}_{\infty}$. Therefore, by Condition (1), we have $M\sim^{*}\lambda M$.
+  • For $M\in\widetilde{\mathcal{M}}_{\infty}$ and $\lambda>0$, we have $\lambda M\in\widetilde{\mathcal{M}}_{\infty}$. Therefore, by Condition (2), we have $M\sim^{*}\lambda M$.
+  • For $M\in\mathcal{M}_{0}$ and $\lambda>0$, we have $\lambda M\in\mathcal{M}_{0}$. Therefore, by Condition (3), we have $M\sim^{*}\lambda M$.
+  • For $M\in\mathcal{M}_{>0}$ and $\lambda>0$, we have $\lambda M\in\mathcal{M}_{>0}$.
+Since $\succ^{*}$ is induced from the odds ratio, we have $M\sim\lambda M$.
+Therefore, $\succeq^{*}$ satisfies marginal monotonicity.
+We conclude that $\succeq^{*}$ satisfies all the three axioms. However, $\succeq^{*}$ is not induced from the odds ratio. To see this, consider the following two matrices.
+[…]
+Since $M\in\mathcal{M}_{\infty}$ and $M^{\prime}\in\widetilde{\mathcal{M}}_{\infty}$, Condition (1) implies $M\succ^{*}M^{\prime}$. However, according to the odds ratio,
+$I_{O}(M)=+\infty=I_{O}(M^{\prime}).$
+Therefore, $\succeq^{*}$ is not induced from the odds ratio.
+One can also verify this negative result by considering the following two matrices.
+[…]
+Since $M\in\mathcal{M}_{0}$ and $M^{\prime}\in\mathcal{M}_{>0}$, Condition (3) implies $M\succ^{*}M^{\prime}$. However, according to the odds ratio,
+$I_{O}(M)=0<1=I_{O}(M^{\prime}).$
+Therefore, $\succeq^{*}$ is not induced from the odds ratio. We conclude that Theorem CDMZ1 is not correct.
+Chiappori et al.
+(2025) are correct in showing that any preorder satisfying the axioms in Theorem CDMZ1 must coincide with the preorder induced by the odds ratio on the class of positive matching matrices $\mathcal{M}_{>0}$.
+However, this conclusion is not guaranteed on the remaining domain.
+To remedy this problem, we make two modifications.
+First, we strengthen Maximum Homogamy to eliminate the preference gap between matrices in $\mathcal{M}_{\infty}$ and $\widetilde{\mathcal{M}}_{\infty}$.
+  • Maximum Homogamy+. For any $M\in\mathcal{M}$ and $(a,b,c,d)\in\mathcal{M}$ such that $bc=0$, we have $(a,b,c,d)\succeq M$.
+The second modification is to impose Maximum Heterogamy as in Section 4 to gurantee that every matrix in $\mathcal{M}_{0}$ is evaluated as least assortative.
+  • Maximum Heterogamy (for preorders). For each $M\in\mathcal{M}_{>0}$ and $(a,b,c,d)\in\mathcal{M}$ such that $b>0,c>0,a=d=0$ (i.e., $(a,b,c,d)\in\mathcal{M}_{0}$), we have $M\succeq(a,b,c,d)$.
+The following theorem recovers the axiomatization of the odds ratio.
+Theorem 4.
+The odds ratio induces the unique preorder that satisfies marginal monotonicity, marginal independence, maximum homogamy+, and maximum heterogamy.
+Note that maximum homogamy+ and maximum heterogamy uniquely pin down the ordinal ranking between a positive matching matrix and non-positive matching matrix, and between nonpositive matching matrices.
+[^4] Precisely, to prove $M\succ M^{\prime}$ for $M\in\mathcal{M}_{>0}$ and $M^{\prime}\in\mathcal{M}_{0}$, we need marginal monotonicity aside from maximum heterogamy.
+The proof proceeds as follows: by maximum heterogamy, we have $M\succeq M^{\prime}$.
+If $M\sim M^{\prime}$, then letting $M=(a,b,c,d)$ and $M_{\varepsilon}=(a-\varepsilon,b+\varepsilon,c+\varepsilon,d-\varepsilon)\in\mathcal{M}_{>0}$ for a sufficiently small $\varepsilon$, marginal monotonicity implies $M\succ M_{\varepsilon}$.
+But then, we have $M^{\prime}\sim M\succ M_{\varepsilon}$, contradicting maximum heterogamy.
+Since marginal monotonicity and marginal independence imply that the oreorder is induced from the odds ratio on the class of positive matching matrices, Theorem 4 holds.
+5.2 Normalized trace
+Specifically, Chiappori et al. (2025) consider the following index called the normalized trace:
 $I_{tr}(M)=\left\{\begin{array}[]{cl}1&\text{ if }bc=0,\\ \frac{a+d}{a+b+c+d}\in(0,1)&\text{ if }abcd\neq 0,\\ 0&\text{ if }ad=0.\end{array}\right$.
@@
 $\hat{\mathcal{M}}=\{M\in\mathcal{M}\mid bc\neq 0\text{ or }ad\neq 0\}.$
-The following is Theorem 2 of chiappori2025changes.
+The following is Theorem 2 of Chiappori et al. (2025).
 Theorem CDMZ2.
@@
 The last two axioms in the statement are defined as follows:
-  • Maximum Homogamy. For each $M\in\hat{\mathcal{M}}$ and $(a,b,c,d)\in\hat{\mathcal{M}}$ such that $ad>0$ and $bc=0$, we have $I(a,b,c,d)\geq I(M)$.
+  • Maximum Homogamy (for indices). For each $M\in\hat{\mathcal{M}}$ and $(a,b,c,d)\in\hat{\mathcal{M}}$ such that $ad>0$ and $bc=0$, we have $I(a,b,c,d)\geq I(M)$.
   • Population Decomposability. For each pair of positive matchings $M,M^{\prime}\in\hat{\mathcal{M}}$,
@@
 One can verify that this modified index satisfies the three invariance axioms, marginal monotonicity, maximum homogamy, and population decomposability.
-[^4] This modified index also satisfies other two monotonicity axioms in chiappori2025changes, namely diagonal monotonicity and off-diagonal monotonicity.
+[^5] This modified index also satisfies other two monotonicity axioms in Chiappori et al.
+(2025), namely diagonal monotonicity and off-diagonal monotonicity.
 However, this modified index is not a positive affine transformation of the normalized trace.
@@
 By adopting a technique similar to that developed in Section 4, we can recover the axiomatization of the normalized trace. We will provide details in an updated version of this note.
+6 Multi-type odds ratio
+In this section we offer a multi-type generalization of the odds ratio.
+The set of types is given by $T=\{1,\dots,n\}$.
+Let $\mathcal{T}=T\times T$ denote the set of type profiles of men and women with generic element $\tau\in\mathcal{T}$.
+A matching $M=(m_{\tau})$ is a matrix in $\mathbb{R}_{\geq 0}^{\mathcal{T}}\setminus\{\textbf{0}\}$ with rows corresponding to men types and columns corresponding to women types.
+For each type profile $\tau=(t,t^{\prime})\in\mathcal{T}$, $m_{\tau}$ denotes the number of observed matched pairs of men of type $t$ and women of type $t^{\prime}$.
+Let $\mathcal{M}_{>0}$ denote the set of all positive matching matrices, i.e., all entries are positive.
+Throughout this section, we define an index as a function $I:\mathcal{M}_{>0}\rightarrow\mathbb{R}$.
+We often write $M\geqslant M^{\prime}$ to denote $I(M)\geq I\left(M^{\prime}\right)$.
+For $M=(m_{\tau})\in\mathcal{M}_{>0}$, $\theta\in\mathcal{T}$, and $m^{\prime}_{\theta}\in\mathbb{R}_{>0}$, let $(M_{-\theta},m^{\prime}_{\theta})\in\mathcal{M}_{>0}$ denote the matching obtained from $M$ by replacing the $\theta$-th entry with $m^{\prime}_{\theta}$, while keeping all other entries the same.
+  • Cell Scale Independence: Let $M=(m_{\tau})\in\mathcal{M}_{>0}$ and $\tilde{M}=(\tilde{m}_{\tau})\in\mathcal{M}_{>0}$. Then, for each $\theta=(t,t^{\prime})\in\mathcal{T}$ and $\lambda>0$,
+    $I(M)>I(\tilde{M})\Longleftrightarrow I(M_{-\theta},\lambda m_{\theta})>I(\tilde{M}_{-\theta},\lambda\tilde{m}_{\theta}).$
+This axiom was implicitly considered by Osborne (1976) in the context of axiomatizing social welfare functions.
+[^6] The main objective of Osborne (1976) is to establish the nonexistence of a social welfare function satisfying a set of axioms, none of which corresponds to Cell Scale Independence.
+However, this condition appears in the Lemma of the paper (p.1007).
+It states that if, in both matchings, the count in any single cell (i.e., a given man–woman type pair) is multiplied by the same positive factor, then the index preserves the ranking of assortativeness between the two matchings.
+This axiom makes sense if we posit that the key determinant of assortativeness is the relative number of matched pairs (compared across types), because after the change, both matrices receive the same proportional change in a particular type pair.
+Note that the odds ratio in Section 5.1 assigns the number $a^{1}\cdot b^{-1}\cdot c^{-1}\cdot d^{1}$ to positive matrices $(a,b,c,d)\in\mathcal{M}_{>0}$.
+Below we generalize this index to multi-type markets.
+For an index $I:\mathcal{M}_{>0}\rightarrow\mathbb{R}$, we define Scale Invariance in the same way as in Section 2.
+An index $I$ satisfies Continuity if $I$ is a continuous function.
+We say that $Z=(z_{\tau})\in\mathbb{R}^{\mathcal{T}}$ is a matrix with zero total sum if $\sum_{\tau\in\mathcal{T}}z_{\tau}=0$.
+Theorem 6.
+$I:\mathcal{M}_{>0}\rightarrow\mathbb{R}$ satisfies Scale Invariance, Continuity, and Cell Scale Independence if and only if there are increasing $\xi:\mathbb{R}\rightarrow\mathbb{R}$ and a matrix $Z=(z_{\tau})\in\mathbb{R}^{\mathcal{T}}$ with zero total sum such that $I(M)=\xi\left(\prod_{\tau\in\mathcal{T}}m_{\tau}^{z_{\tau}}\right)$ for all $M=(m_{\tau})\in\mathcal{M}_{>0}$.
+Proof.
+Proof of the “if” direction: Suppose that there are increasing $\xi:\mathbb{R}\rightarrow\mathbb{R}$ and a matrix $Z=(z_{\tau})\in\mathbb{R}^{\mathcal{T}}$ with zero total sum such that $I$ is given as stated.
+Then, it is clear that $I$ satisfies Continuity.
+We show that $I$ satisfies Scale Invariance.
+Consider $M=(m_{\tau})\in\mathcal{M}_{>0}$ and $\lambda>0$.
+Our goal is to show that $I(\lambda M)=I(M)$.
+Since both $\xi$ and $\log$ are strictly increasing functions, it suffices to show that
+$\log\left(\prod_{\tau\in\mathcal{T}}m_{\tau}^{z_{\tau}}\right)=\log\left(\prod_{\tau\in\mathcal{T}}(\lambda m_{\tau})^{z_{\tau}}\right).$ (19)
+Rewriting the right-hand side,
+$\log\left(\prod_{\tau\in\mathcal{T}}(\lambda m_{\tau})^{z_{\tau}}\right)$ $=\sum_{\tau\in\mathcal{T}}z_{\tau}\log\lambda m_{\tau}=\sum_{\tau\in\mathcal{T}}z_{\tau}\log\lambda+\sum_{\tau\in\mathcal{T}}z_{\tau}\log m_{\tau}=\sum_{\tau\in\mathcal{T}}z_{\tau}\log m_{\tau},$
+where the last equality follows from the assumption that $Z$ is a matrix with zero total sum, i.e., $\sum_{\tau\in\mathcal{T}}z_{\tau}=0$.
+Since the right-most term is equal to the left-hand side of (19), the desired equation holds.
+Next, we show that $I$ satisfies Cell Scale Independence. Consider $M=(m_{\tau})\in\mathcal{M}_{>0}$, $\tilde{M}=(\tilde{m}_{\tau})\in\mathcal{M}_{>0}$, $\theta\in\mathcal{T}$ and $\lambda>0$.
+$I(M~)\displaystyle$ $I(\tilde{M})$ $\Longleftrightarrow\xi\left(\prod_{\tau\in\mathcal{T}}m_{\tau}^{z_{\tau}}\right)>\xi\left(\prod_{\tau\in\mathcal{T}}\tilde{m}_{\tau}^{z_{\tau}}\right)$
+$\Longleftrightarrow\prod_{\tau\in\mathcal{T}}m_{\tau}^{z_{\tau}}>\prod_{\tau\in\mathcal{T}}\tilde{m}_{\tau}^{z_{\tau}}$
+$\Longleftrightarrow\log\left(\prod_{\tau\in\mathcal{T}}m_{\tau}^{z_{\tau}}\right)>\log\left(\prod_{\tau\in\mathcal{T}}\tilde{m}_{\tau}^{z_{\tau}}\right)$
+$\Longleftrightarrow\sum_{\tau\in\mathcal{T}}z_{\tau}\log m_{\tau}>\sum_{\tau\in\mathcal{T}}z_{\tau}\log\tilde{m}_{\tau}$
+$\Longleftrightarrow\sum_{\tau\in\mathcal{T}}z_{\tau}\log m_{\tau}+z_{\theta}\cdot\log\lambda>\sum_{\tau\in\mathcal{T}}z_{\tau}\log\tilde{m}_{\tau}+z_{\theta}\cdot\log\lambda$
+$\Longleftrightarrow\sum_{\tau\in\mathcal{T}\setminus\{\theta\}}z_{\tau}\log m_{\tau}+z_{\theta}\cdot\log(\lambda\cdot m_{\theta})>\sum_{\tau\in\mathcal{T}\setminus\{\theta\}}z_{\tau}\log\tilde{m}_{\tau}+z_{\theta}\cdot\log(\lambda\cdot\tilde{m}_{\theta})$
+$\Longleftrightarrow\log\left(\prod_{\tau\in\mathcal{T}\setminus\{\theta\}}m_{\tau}^{z_{\tau}}\cdot(\lambda\cdot m_{\theta})^{z_{\theta}}\right)>\log\left(\prod_{\tau\in\mathcal{T}\setminus\{\theta\}}\tilde{m}_{\tau}^{z_{\tau}}\cdot(\lambda\cdot\tilde{m}_{\theta})^{z_{\theta}}\right)$
+$\Longleftrightarrow\prod_{\tau\in\mathcal{T}\setminus\{\theta\}}m_{\tau}^{z_{\tau}}\cdot(\lambda\cdot m_{\theta})^{z_{\theta}}>\prod_{\tau\in\mathcal{T}\setminus\{\theta\}}\tilde{m}_{\tau}^{z_{\tau}}\cdot(\lambda\cdot\tilde{m}_{\theta})^{z_{\theta}}$
+$\Longleftrightarrow\xi\left(\prod_{\tau\in\mathcal{T}\setminus\{\theta\}}m_{\tau}^{z_{\tau}}\cdot(\lambda\cdot m_{\theta})^{z_{\theta}}\right)>\xi\left(\prod_{\tau\in\mathcal{T}\setminus\{\theta\}}\tilde{m}_{\tau}^{z_{\tau}}\cdot(\lambda\cdot\tilde{m}_{\theta})^{z_{\theta}}\right)$
+$\Longleftrightarrow I(M_{-\theta},\lambda m_{\theta})>I(\tilde{M}_{-\theta},\lambda\tilde{m}_{\theta}),$
+where the second and the second-to-last equivalences follow from the fact that $\xi$ is an increasing function.
+Proof of the “only if” direction: We introduce some preliminaries. Let $\theta\in\mathcal{T}$. We say that $I$ is nondecreasing in $\theta$ if
+$I(M)\leq I(M_{-\theta},m^{\prime}_{\theta})\text{ for all }M=(m_{\tau})\in\mathcal{M}_{>0}\text{ and }m^{\prime}_{\theta}\in\mathbb{R}_{>0}\text{ with }m^{\prime}_{\theta}\geq m_{\theta}.$
+We say that $I$ is nonincreasing in $\theta$ if
+$I(M)\geq I(M_{-\theta},m^{\prime}_{\theta})\text{ for all }M=(m_{\tau})\in\mathcal{M}_{>0}\text{ and }m^{\prime}_{\theta}\in\mathbb{R}_{>0}\text{ with }m^{\prime}_{\theta}\geq m_{\theta}.$
+Let 1l denote the matching in which all the entries are equal to 1.
+Claim 1.
+Let $I$ be an index that satisfies Continuity and Cell Scale Independence. Then, for each $\theta\in\mathcal{T}$, $I$ is nondecreasing or nonincreasing in $\theta$ (possibly both).
+Proof.
+Fix $\theta\in\mathcal{T}$.
+Suppose, for contradiction, that $I$ is neither nondecreasing nor nonincreasing.
+Then, there exist $M=(m_{\tau})\in\mathcal{M}_{>0}$, $m^{\prime}_{\theta}>m_{\theta}$, and $\tilde{M}=(\tilde{m}_{\tau})\in\mathcal{M}_{>0}$, $\tilde{m}^{\prime}_{\theta}>\tilde{m}_{\theta}$ s.t.
+$I(M)$ $>I(M_{-\theta},m^{\prime}_{\theta}),$ (20)
+[…]
+Consider matchings $M$ and $(M_{-\theta},m^{\prime}_{\theta})$.
+For any $\tau\in\mathcal{T}$ with $\tau\neq\theta$, multiplying the $\tau$-th entry of $M$ and of $(M_{-\theta},m^{\prime}_{\theta})$ by $1/m_{\tau}$ makes the value of that entry in both matrices equal to $1$.
+By Cell Scale Independence, (20) is equivalent to
+$I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},m_{\theta})>I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},m^{\prime}_{\theta}).$ (22)
+Similarly, (21) is equivalent to
+$I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},\tilde{m}_{\theta})<I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},\tilde{m}^{\prime}_{\theta}).$ (23)
+We define $f:\mathbb{R}_{>0}\rightarrow\mathbb{R}$ by
+$f(x_{\theta})=I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},x_{\theta})\text{ for all }x_{\theta}\in\mathbb{R}_{>0}.$
+Then, (22) and (23) are rewritten as
+$f(m_{\theta})>f(m^{\prime}_{\theta}),$ (24)
+$f(\tilde{m}_{\theta})<f(\tilde{m}^{\prime}_{\theta}).$ (25)
+We define
+$m^{*}_{\theta}:=\sup\{x_{\theta}\in[m_{\theta},m^{\prime}_{\theta}]\mid f(x_{\theta})\geq f(m_{\theta})\}.$
+Since $m^{*}_{\theta}$ is defined by the supremum, there exists a sequence $\{x^{k}_{\theta}\}_{k=1}^{\infty}\subseteq[m_{\theta},m^{\prime}_{\theta}]$ such that $f(x^{k}_{\theta})\geq f(m_{\theta})$ for all $k=1,2,\dots$, and $x^{k}_{\theta}\rightarrow m^{*}_{\theta}$ as $k\rightarrow\infty$.
+By Continuity,
+$\lim_{k\rightarrow\infty}f(x_{\theta}^{k})=f(m^{*}_{\theta})\geq f(m_{\theta}).$
+This ineuqality together with (24) implies $m^{*}_{\theta}<m^{\prime}_{\theta}$.
+Since $m^{*}_{\theta}$ is the supremum of points with a function value no less than $f(m_{\theta})$, any point $x_{\theta}>m^{*}_{\theta}$ attains a higher functioin value than $f(m_{\theta})$.
+Therefore, we have
+$f(x_{\theta})<f(m_{\theta})\leq f(m^{*}_{\theta})\text{ for all }x_{\theta}\in(m^{*}_{\theta},m^{\prime}_{\theta}]$ (26)
+So far we have considered matchings $M$ and $(M_{-\theta},m^{\prime}_{\theta})$. Analogously we consider matchings $\tilde{M}$ and $(\tilde{M}_{-\theta},\tilde{m}^{\prime}_{\theta})$. We define
+$\tilde{m}^{*}_{\theta}:=\sup\{x_{\theta}\in[\tilde{m}_{\theta},\tilde{m}^{\prime}_{\theta}]\mid f(x_{\theta})\leq f(\tilde{m}_{\theta})\}.$
+By following the same argument used in the previous paragraph (while using (25) instead of (24)), we obtain $\tilde{m}^{*}_{\theta}<\tilde{m}^{\prime}_{\theta}$ and
+$f(x_{\theta})>f(\tilde{m}_{\theta})\geq f(\tilde{m}^{*}_{\theta})\text{ for all }x_{\theta}\in(\tilde{m}^{*}_{\theta},\tilde{m}^{\prime}_{\theta}].$ (27)
+We define
+$\alpha:=\min\Bigl\{\frac{m^{\prime}_{\theta}}{m^{*}_{\theta}},\frac{\tilde{m}^{\prime}_{\theta}}{\tilde{m}^{*}_{\theta}}\Bigr\}.$
+Since $m^{*}_{\theta}<m^{\prime}_{\theta}$ and $\tilde{m}^{*}_{\theta}<\tilde{m}^{\prime}_{\theta}$, we have $\alpha>1$. The following inequalities hold:
+$m^{*}_{\theta}<\alpha\cdot m^{*}_{\theta}\leq\frac{m^{\prime}_{\theta}}{m^{*}_{\theta}}\cdot m^{*}_{\theta}=m^{\prime}_{\theta},$
+where the first inequality follows from $m^{*}_{\theta}>0$ and $\alpha>1$, and the second inequality follows from the definition of $\alpha$. Similarly,
+$\tilde{m}^{*}_{\theta}<\alpha\cdot\tilde{m}^{*}_{\theta}\leq\frac{\tilde{m}^{\prime}_{\theta}}{\tilde{m}^{*}_{\theta}}\cdot\tilde{m}^{*}_{\theta}=\tilde{m}^{\prime}_{\theta},$
+By (26) for $x_{\theta}\leftarrow\alpha\cdot m^{*}_{\theta}$ and (27) for $x_{\theta}\leftarrow\alpha\cdot\tilde{m}^{*}_{\theta}$, we have
+$f(m^{*}_{\theta})>f(\alpha\cdot m^{*}_{\theta}),$
+$f(\tilde{m}^{*}_{\theta})<f(\alpha\cdot\tilde{m}^{*}_{\theta}).$
+By the definition of $f$,
+$I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},m^{*}_{\theta})>I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},\alpha\cdot m^{*}_{\theta}),$
+$I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},\tilde{m}^{*}_{\theta})<I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},\alpha\cdot\tilde{m}^{*}_{\theta}).$
+The former inequality means that the matching $(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},m^{*}_{\theta})$ is more assortative than the matching $(\textup{\mbox{1}\hskip-2.5pt\mbox{l}}_{-\theta},\alpha\cdot m^{*}_{\theta})$.
+Multiplying the $\theta$-th entry of both matrices by $\frac{\tilde{m}^{*}_{\theta}}{m^{*}_{\theta}}$, we obtain two matchings that appear in the latter inequality, where assortativeness is reversed.
+We obtain a contradiction to Cell Scale Independence.
+∎
+We define
+$\mathcal{T}_{-}=\{\theta\in\mathcal{T}\mid\text{$I$ is not nondecreasing in $\theta$}\}.$
+For $M=(m_{\tau})\in\mathcal{M}_{>0}$, we define $M^{\text{inv}(\mathcal{T}_{-})}$ as follows: for each $\tau\in\mathcal{T}$,
+$(M^{\text{inv}(\mathcal{T}_{-})})_{\tau}=\begin{cases}1/m_{\tau}&\text{ if }\tau\in\mathcal{T}_{-},\\ m_{\tau}&\text{ otherwise. }\end{cases}$
+We define a new index $I^{\text{inv}(\mathcal{T}_{<0})}$ as follows:
+$I^{\text{inv}(\mathcal{T}_{-})}(M)=I(M^{\text{inv}(\mathcal{T}_{-})})\text{ for all }M\in\mathcal{M}_{>0}.$
+Claim 2.
+For each $\theta\in\mathcal{T}$, $I^{\text{inv}(\mathcal{T}_{-})}$ is nondecreasing in $\theta$.
+Proof.
+Consider $\theta\in\mathcal{T}$.
+If $I$ is nondecreasing in $\theta$, then it is clear that $I^{\text{inv}(\mathcal{T}_{-})}$ is also nondecreasing in $\theta$.
+Suppose that $I$ is not nondecreasing in $\theta$, i.e., $\theta\in\mathcal{T}_{-}$.
+By Claim 1, $I$ is nonincreasing in $\theta$.
+Then, for any $M=(m_{\tau})\in\mathcal{M}_{>0}$ and $m^{\prime}_{\theta}\in\mathbb{R}_{>0}$ with $m^{\prime}_{\theta}\geq m_{\theta}$,
+$I^{\text{inv}(\mathcal{T}_{-})}(M)$ $=I(M^{\text{inv}(\mathcal{T}_{-})})$
+$=I\Bigl(M^{\text{inv}(\mathcal{T}_{-})}_{-\theta},\frac{1}{m_{\theta}}\Bigr)$
+$\leq I\Bigl(M^{\text{inv}(\mathcal{T}_{-})}_{-\theta},\frac{1}{m^{\prime}_{\theta}}\Bigr)$
+$=I\Bigl((M_{-\theta},m^{\prime}_{\theta})^{\text{inv}(\mathcal{T}_{-1})}\Bigr)$
+$=I^{\text{inv}(\mathcal{T}_{-})}(M_{-\theta},m^{\prime}_{\theta}),$
+where the second and the second-to-last equalities follow from $\theta\in\mathcal{T}_{-}$ and the inequality follows from the fact that $I$ is nonincreasing in $\theta$. ∎
+Claim 3.
+$I^{\text{inv}(\mathcal{T}_{-})}$ satisfies Cell Scale Independence.
+Proof.
+Consider $M=(m_{\tau})\in\mathcal{M}_{>0}$ and $\tilde{M}=(\tilde{m}_{\tau})\in\mathcal{M}_{>0}$, $\theta\in\mathcal{T}$, and $\lambda>0$.
+If $\theta\notin\mathcal{T}_{-}$, then the claim immediately follows from Cell Scale Independence of $I$.
+Suppose that $\theta\in\mathcal{T}_{-}$.
+[…]
+$\Longleftrightarrow I\left(M^{\text{inv}(\mathcal{T}_{-})}_{-\theta},\frac{1}{\lambda}\cdot\frac{1}{m_{\theta}}\right)>I\left(\tilde{M}^{\text{inv}(\mathcal{T}_{-})}_{-\theta},\frac{1}{\lambda}\cdot\frac{1}{\tilde{m}_{\theta}}\right)$
+$\Longleftrightarrow I^{\text{inv}(\mathcal{T}_{-})}\left(M_{-\theta},\lambda m_{\theta}\right)>I^{\text{inv}(\mathcal{T}_{-})}(\tilde{M}_{-\theta},\lambda\tilde{m}_{\theta}),$
+where the second equivalence follows from Scale Cell Independence of $I$. ∎
+We make use of the following characterization of nondecreasing and cell-scale independent functions.
+Proposition 1 (Lemma of Osborne (1976)).
+An index $I$ is nondecreasing in each $\theta\in\mathcal{T}$ and satisfies Cell Scale Independence if and only if there are increasing $\xi:\mathbb{R}\rightarrow\mathbb{R}$ and a matrix $Z=(z_{\tau})\in\mathbb{R}^{\mathcal{T}}_{\geq 0}$ such that $I(M)=\xi\left(\prod_{\tau\in\mathcal{T}}m_{\tau}^{z_{\tau}}\right)$ for all $M=(m_{\tau})\in\mathcal{M}_{>0}$.
+Combining Claims 1 and 2 with Proposition 1, there are increasing $\xi:\mathbb{R}\rightarrow\mathbb{R}$ and a matrix $Z=(z_{\tau})\in\mathbb{R}^{\mathcal{T}}_{\geq 0}$ such that $I^{\text{inv}(\mathcal{T}_{-})}(M)=\xi\left(\prod_{\tau\in\mathcal{T}}m_{\tau}^{z_{\tau}}\right)$ for all $M=(m_{\tau})\in\mathcal{M}_{>0}$.
+Then, for each $M=(m_{\tau})\in\mathcal{M}_{>0}$,
+$I(M)$ $=I^{\text{inv}(\mathcal{T}_{-})}(M^{\text{inv}(\mathcal{T}_{-})})$
+$=\xi\left(\prod_{\tau\in\mathcal{T}\setminus\mathcal{T}_{-}}m_{\tau}^{z_{\tau}}\prod_{\theta\in\mathcal{T}_{-}}\left(\frac{1}{m_{\theta}}\right)^{z_{\theta}}\right)$
+$=\xi\left(\prod_{\tau\in\mathcal{T}\setminus\mathcal{T}_{-}}m_{\tau}^{z_{\tau}}\prod_{\theta\in\mathcal{T}_{-}}m_{\theta}^{-z_{\theta}}\right)$
+Therefore, the desired condition holds for the matrix $Z^{\text{inv}(\mathcal{T}_{-})}$.
+Finally, we show that $Z^{\text{inv}(\mathcal{T}_{-})}$ has zero total sum.
+Let $\bar{z}$ denote the sum of the entries of this matrix.
+For matrices 1l and $2\cdot\textup{\mbox{1}\hskip-2.5pt\mbox{l}}$,
+$I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}})=\xi(1),\>I(2\cdot\textup{\mbox{1}\hskip-2.5pt\mbox{l}})=\xi(2^{\bar{z}}).$
+By Scale Invariance, $I(\textup{\mbox{1}\hskip-2.5pt\mbox{l}})=I(2\cdot\textup{\mbox{1}\hskip-2.5pt\mbox{l}})$.
+Since $\xi(\cdot)$ is an increasing function, we have $1=2^{\bar{z}}$.
+It follows that the sum $\bar{z}$ must be equal to 0, as desired.
+∎
+One can consider the following weakening of Cell Scale Independence.
+  • Cell Scale Independence-: Let $M=(m_{\tau})\in\mathcal{M}_{>0}$ and $\tilde{M}=(\tilde{m}_{\tau})\in\mathcal{M}_{>0}$.
+Then, for each $\theta=(t,t^{\prime})\in\mathcal{T}$ with $t\neq t^{\prime}$ and $\lambda>0$,
+    $I(M)>I(\tilde{M})\Longleftrightarrow I(M_{-\theta},\lambda m_{\theta})>I(\tilde{M}_{-\theta},\lambda\tilde{m}_{\theta}).$
+This axiom states that, if the ratio of the number of different-type pairs to that of same-type pairs is scaled by the same factor in $M$ and $\tilde{M}$, then the relative assortativeness between $M$ and $\tilde{M}$ should remain unchanged.
+This axiom, together with a multi-type version of Marginal Independence (see Section 5.1), yields an axiomatization of the same form as in Theorem 6, although the class of matrices $Z$ is more restricted.
+The class of matrices $Z$ can also be restricted by introducing multi-type versions of Type Invariance and Marginal Monotonicity (see Section 2).
+We discuss these issues in an updated version of this note.